# NVS-SQA: Exploring Self-Supervised Quality Representation Learning for Neurally Synthesized Scenes without References

### v1

@@ IV-A Datasets and evaluation protocol
Labeled datasets for evaluation. The evaluation of NVS-SQA leveraged three distinct labeled NSS datasets—Lab [5], LLFF [10], and Fieldwork [5]—each presenting unique challenges. The Lab dataset features six real scenes captured with a 2D gantry system in a laboratory setting, offering a uniform grid of training views and reference videos ranging from 300 to 500 frames for a thorough quality assessment. In contrast, the LLFF dataset comprises eight real scenes captured with a handheld cellphone, providing a sparse selection of 20-30 test views per scene, with positional data computed via COLMAP [46]. The Fieldwork dataset includes nine real scenes from diverse environments, such as outdoor urban landscapes and indoor museum settings, characterized by intricate backgrounds and variable lighting conditions, with reference videos typically containing around 120 frames. To synthesize NSS for each scene, ten NVS methods were utilized, spanning a variety of models with explicit and implicit geometric representations, rendering models, and optimization strategies, ensuring an unbiased evaluation of the framework’s generalizability. Moreover, methods such as NeRF [1], Mip-NeRF [40], DVGO [41], Plenoxels [42], NeX [44], LFNR [45], IBRNet [59], and GNT [60] were involved, including both cross-scene (GNT-C and IBRNet-C) and scene-specific (GNT-S and IBRNet-S) models. The human perceptual labels were collected in a comparison-based manner and in format of Just-Objectionable-Difference (JOD) units [5]. It is important to highlight that the JOD scale’s utility and the comparative analysis it facilitates are particularly pronounced when conducted within the context of individual scenes [5].
@@
Evaluation protocol. We largely adhere to the evaluation protocol from [29, 30, 16], but in a more stringent manner. Our evaluation consists of two primary objectives: first, to determine if the proposed method can simultaneously adapt to three different datasets; second, to test the method’s ability to perform cross-dataset validation. Following the self-supervised learning phase, the model’s weights are frozen, and the model is applied to the three evaluation datasets to generate quality representations. A one-time linear regression is performed on a randomly sampled half of the integrated dataset to align these quality representations with human scores. The model’s performance is then evaluated on the remaining half of the dataset to assess the generalizability of the NVS-SQA framework across the integrated datasets. This approach is designed to thoroughly evaluate the framework’s capability to adapt and perform effectively amidst the varied complexities posed by the Lab, LLFF, and Fieldwork datasets simultaneously. Additionally, we conduct cross-dataset experiments to further assess the adaptability of the proposed method across various datasets.
@@
Training setup. The training regimen for the model was executed using the ADAM optimizer [64], spanning 200 epochs with a batch size of 16. Consequently, once the training on an unlabeled dataset is complete, the model is adept at being applied proficiently to novel scenes across varied datasets after minor linear regression. The computational experiments underpinning this research were performed on a high-specification desktop equipped with an AMD 5950X processor, an RTX 4090 GPU, and 128GB of RAM, running on the Windows 10 operating system. The implementation was carried out in PyTorch [65].
@@
TV +$0.378$ ($0.64$) +$0.423$ ($0.58$) +$0.311$ ($0.62$) +$0.087$ ($0.65$) +$0.050$ ($0.61$) +$0.075$ ($0.52$) +$0.200$ ($0.48$) +$0.136$ ($0.30$) +$\underline{0.229}$ ($0.36$)
BRISQUE +$0.089$ ($0.58$) +$0.152$ ($0.58$) +$0.067$ ($0.53$) $-0.037$ ($0.54$) $-0.103$ ($0.45$) $-0.050$ ($0.44$) +$\underline{0.214}$ ($0.83$) +$0.204$ ($0.72$) +$0.171$ ($0.70$)
NIQE +$0.467$ ($0.63$) +$0.331$ ($0.66$) +$0.400$ ($0.56$) +$0.025$ ($0.50$) $-0.077$ ($0.43$) +$0.025$ ($0.42$) $-0.357$ ($0.16$) $-0.329$ ($0.12$) $-0.314$ ($0.17$)
PIQE +$0.079$ ($0.49$) $-0.079$ ($0.55$) +$0.273$ ($0.57$) +$0.047$ ($0.56$) $-0.079$ ($0.47$) $-0.021$ ($0.44$) $-0.364$ ($0.65$) $-0.060$ ($0.70$) $-0.302$ ($0.49$)
CLIP-IQA +$0.233$ ($0.63$) +$0.178$ ($0.57$) +$0.200$ ($0.54$) +$0.025$ ($0.47$) $-0.046$ ($0.36$) +$0.025$ ($0.34$) $-0.057$ ($0.50$) $-0.240$ ($0.47$) $-0.086$ ($0.36$)
CONTRIQUE +$\underline{0.689}$ ($0.28$) +$\underline{0.759}$ ($0.29$) +$\underline{0.622}$ ($0.29$) +$0.350$ ($0.40$) +$0.400$ ($0.50$) +$0.275$ ($0.33$) +$0.086$ ($0.43$) +$0.200$ ($0.53$) +$0.057$ ($0.36$)
Re-IQA +$0.589$ ($0.53$) +$0.585$ ($0.48$) +$0.489$ ($0.44$) +$0.062$ ($0.71$) $-0.018$ ($0.71$) +$0.025$ ($0.61$) +$0.143$ ($0.16$) +$\underline{0.213}$ ($0.30$) +$0.200$ ($0.10$)
VIIDEO +$0.022$ ($0.37$) +$0.070$ ($0.44$) +$0.022$ ($0.28$) $-0.050$ ($0.49$) +$0.002$ ($0.49$) $-0.000$ ($0.39$) +$0.000$ ($0.27$) +$0.061$ ($0.47$) $-0.029$ ($0.30$)
Video-BlIINDS +$0.189$ ($0.39$) +$0.148$ ($0.41$) +$0.156$ ($0.31$) +$0.162$ ($0.44$) +$0.051$ ($0.40$) +$0.175$ ($0.38$) $-0.314$ ($0.33$) $-0.113$ ($0.42$) $-0.286$ ($0.33$)
FAST-VQA +$0.167$ ($0.59$) +$0.106$ ($0.59$) +$0.122$ ($0.46$) +$0.112$ ($0.55$) +$0.255$ ($0.55$) +$0.185$ ($0.41$) $-0.171$ ($0.65$) $-0.191$ ($0.63$) $-0.186$ ($0.52$)
FasterVQA +$0.186$ ($0.51$) +$0.245$ ($0.55$) +$0.172$ ($0.47$) +$0.162$ ($0.58$) +$0.176$ ($0.49$) +$0.125$ ($0.48$) $-0.129$ ($0.46$) $-0.175$ ($0.44$) $-0.186$ ($0.38$)
DOVER +$0.200$ ($0.56$) +$0.267$ ($0.65$) +$0.267$ ($0.50$) +$0.150$ ($0.61$) +$0.153$ ($0.47$) +$0.125$ ($0.48$) $-0.129$ ($0.26$) $-0.209$ ($0.18$) $-0.214$ ($0.22$)
DOVER-Mobile +$0.344$ ($0.60$) +$0.341$ ($0.61$) +$0.311$ ($0.47$) +$0.200$ ($0.47$) +$0.254$ ($0.36$) +$0.125$ ($0.40$) $-0.071$ ($0.47$) $-0.097$ ($0.32$) $-0.143$ ($0.46$)
NR-LFQA +$0.207$ ($0.52$) +$0.157$ ($0.56$) +$0.149$ ($0.42$) +$0.088$ ($0.47$) +$0.106$ ($0.34$) +$0.114$ ($0.37$) $-0.094$ ($0.59$) $-0.067$ ($0.57$) $-0.039$ ($0.51$)
Tensor-NLFQ +$0.282$ ($0.42$) +$0.290$ ($0.49$) +$0.202$ ($0.43$) +$0.223$ ($0.38$) +$0.210$ ($0.36$) +$0.125$ ($0.29$) $-0.061$ ($0.64$) $-0.064$ ($0.57$) $-0.072$ ($0.55$)
@@
LFACon +$0.333$ ($0.44$) +$0.310$ ($0.47$) +$0.242$ ($0.33$) +$\underline{0.412}$ ($0.48$) +$\underline{0.395}$ ($0.31$) +$\underline{0.325}$ ($0.36$) +$0.157$ ($0.60$) +$0.132$ ($0.53$) +$0.086$ ($0.51$)
Proposed +$\boldsymbol{0.911}$ ($0.08$) +$\boldsymbol{0.883}$ ($0.11$) +$\boldsymbol{0.822}$ ($0.16$) +$\boldsymbol{0.700}$ ($0.28$) +$\boldsymbol{0.644}$ ($0.39$) +$\boldsymbol{0.625}$ ($0.19$) +$\boldsymbol{0.700}$ ($0.04$) +$\boldsymbol{0.678}$ ($0.10$) +$\boldsymbol{0.571}$ ($0.09$)
@@
Figure 7: How often does each quality estimation method align with ground-truth perceptual labels (indicated by /’-shaped lines)? We compare our proposed NVS-SQA against three competitive representatives of no-reference quality assessment: CONTRIQUE (IQA), DOVER (VQA), and LFACon (LFIQA), evaluated on the Fieldwork, LLFF, and Lab datasets. Each subfigure shows scatter plots where different symbols and colors represent individual scenes, with a linear regression line for each. The subfigure titles (e.g., [CONTRIQUE] on LLFF (2/8)‘) indicate how many of these lines have a positive slope, reflecting stronger agreement between the estimated and ground-truth scores. A higher number of positively sloped lines signifies more robust alignment of the method’s predictions with perceptual labels across multiple scenes.
@@ IV-B Comparison with no-reference quality assessment methods
Visualiazation of correlation with perceptual labels. Figure 7 illustrates how well the predicted quality assessments correlate with ground-truth perceptual labels. Specifically, each positively sloped line indicates a strong alignment between the predicted quality and human judgments (further explained in the caption). As shown in the figure, NVS-SQA achieves the highest number of positively sloped lines in every dataset, reaching perfect alignment in Fieldwork (9 of 9 lines) and top performance in both LLFF (7 of 8) and Lab (7 of 7). This result underscores the robust effectiveness of NVS-SQA across a diverse set of scenes and highlights its superior alignment with human perception compared to existing approaches.
@@
TV +$0.344$ ($0.58$) +$0.398$ ($0.58$) +$0.244$ ($0.54$) $-0.050$ ($0.50$) $-0.032$ ($0.53$) +$0.025$ ($0.42$) +$0.143$ ($0.51$) +$0.266$ ($0.27$) +$0.171$ ($0.46$)
BRISQUE $-0.044$ ($0.59$) $-0.004$ ($0.58$) $-0.067$ ($0.51$) $-0.237$ ($0.50$) $-0.289$ ($0.52$) $-0.175$ ($0.42$) +$\underline{0.214}$ ($0.42$) +$0.279$ ($0.43$) +$0.171$ ($0.35$)
NIQE +$0.344$ ($0.37$) +$0.337$ ($0.40$) +$0.311$ ($0.33$) $-0.212$ ($0.58$) $-0.138$ ($0.55$) $-0.175$ ($0.50$) $-0.243$ ($0.22$) $-0.188$ ($0.41$) $-0.171$ ($0.22$)
PIQE $-0.092$ ($0.51$) $-0.012$ ($0.58$) $-0.211$ ($0.56$) $-0.175$ ($0.55$) $-0.173$ ($0.51$) $-0.171$ ($0.54$) $-0.169$ ($0.30$) $-0.237$ ($0.25$) $-0.295$ ($0.21$)
CLIP-IQA +$0.444$ ($0.34$) +$0.525$ ($0.32$) +$0.356$ ($0.25$) +$0.175$ ($0.60$) +$0.264$ ($0.67$) +$0.100$ ($0.49$) $-0.057$ ($0.25$) +$0.020$ ($0.26$) $-0.029$ ($0.22$)
CONTRIQUE +$\underline{0.622}$ ($0.38$) +$\underline{0.626}$ ($0.31$) +$\underline{0.522}$ ($0.42$) $-0.013$ ($0.34$) +$0.050$ ($0.48$) $-0.075$ ($0.20$) $-0.143$ ($0.33$) $-0.102$ ($0.56$) $-0.114$ ($0.32$)

### v3

@@ III-E AdaptiSceneNet
III-F Self-supervised training procedure
To complement the contrastive pair generation described in Algorithm 1, we present the self-supervised training procedure in Algorithm 2 for NVS-SQA, which includes batch-wise pair generation, multi-branch loss computation, and model updates using the AdaptiSceneNet backbone.
Algorithm 2 Self-Supervised Training Procedure
1:Unlabeled dataset $D$ of NSS, number of training epochs $E$, batch size $B$, learning rate $\eta$
2:Trained quality representation model $H$
3:Initialize model weights for AdaptiSceneNet
4:for $\text{epoch}\leftarrow 1$ to $E$ do
5:  $P\leftarrow\text{GenerateContrastivePairs}(D,\text{batch size}=B)$ $\triangleright$ Using Algorithm 1
6:  for each $(s_{1},s_{2})\in P$ do
7:    Compute quality representations $H(s_{1}),H(s_{2})$
8:    for each branch do $\psi\in\{\text{IQA},\text{VQA},\text{REP}\}$
9:     Compute branch loss $\ell_{\psi}$ using Eq. 7
10:    end for
11:    Aggregate total loss
12:    Backpropagate and update network weights
13:  end for
14:end for
15:return $H$
Time complexity analysis. Let $N$ be the number of training scenes, $V$ the number of views per scene, $R\times C$ the spatial resolution of each view, $d$ the feature dimension, and $L$ the number of transformer layers. The contrastive pair generation step (Algorithm 1) has complexity $\mathcal{O}(N)$, as each scene is processed independently. For each training step, the AdaptiSceneNet backbone first applies viewwise convolutions with complexity $\mathcal{O}(B\cdot V\cdot R\cdot C)$ for a batch size $B$. It then performs transformer-based angular fusion with self-attention, which has complexity $\mathcal{O}(B\cdot L\cdot V^{2}\cdot d)$. MLP heads have negligible cost in comparison. Hence, the total per-batch time complexity is $\mathcal{O}(B\cdot V\cdot R\cdot C+B\cdot L\cdot V^{2}\cdot d)$. Despite the quadratic scaling in $V$, the model training remains computationally practical. On an RTX 4090 GPU, the average training time per NSS pair is approximately 8 seconds. During inference, our method requires only 0.692 seconds per scene on average. Further details about computational cost analysis are provided in Section IV-G.
This study involved human subjects for the collection of perceptual ratings in the newly introduced dataset (as described in Section IV-H). All experimental procedures and protocols were approved by the Research Ethics Review Committee of BTBU. Informed consent was obtained from all participants prior to data collection.
@@ IV-A Datasets and evaluation protocol
Training data diversity and pair generation scale. In addition to this scalable supervision, the scenes themselves were carefully selected to maximize real-world diversity. Specifically, these scenes were obtained using different devices such as mobile phones and mirrorless cameras with fisheye lenses [56]. Importantly, this dataset is not limited to canonical forward-facing captures and includes 360-degree scenes with varying levels of quality and structural complexity. Unlike other datasets that focus on a central object or object-centric environments, the scenes span both indoor and outdoor settings, sparse and cluttered layouts, and complex motion paths [56]. This ensures that our training data reflect the diversity and imperfections encountered in real-world applications. Finally, the high variability in view counts, spatial resolution, scene structure, and camera trajectories further contributes to the diversity of training signals. These factors, combined with our contrastive pair generation strategy, allow the model to generalize beyond the training distribution, as demonstrated by its consistent performance on multiple unseen datasets.
@@
Labeled datasets for evaluation. The evaluation of NVS-SQA leveraged three distinct labeled NSS datasets—Lab [5], LLFF [10], and Fieldwork [5]—each presenting unique challenges. The Lab dataset features six real scenes captured with a 2D gantry system in a laboratory setting, offering a uniform grid of training views and reference videos ranging from 300 to 500 frames for a thorough quality assessment. In contrast, the LLFF dataset comprises eight real scenes captured with a handheld cellphone, providing a sparse selection of 20-30 test views per scene, with positional data computed via COLMAP [46]. The Fieldwork dataset includes nine real scenes from diverse environments, such as outdoor urban landscapes and indoor museum settings, characterized by intricate backgrounds and variable lighting conditions, with reference videos typically containing around 120 frames. To synthesize NSS for each scene, ten NVS methods were utilized, spanning a variety of models with explicit and implicit geometric representations, rendering models, and optimization strategies, ensuring an unbiased evaluation of the framework’s generalizability. Moreover, methods such as NeRF [1], Mip-NeRF [40], DVGO [41], Plenoxels [42], NeX [44], LFNR [45], IBRNet [59], and GNT [60] were involved, including both cross-scene (GNT-C and IBRNet-C) and scene-specific (GNT-S and IBRNet-S) models. The human perceptual labels were collected in a comparison-based manner and in format of Just-Objectionable-Difference (JOD) units [5]. These JOD scores were obtained from publicly available datasets [5], which were constructed through controlled user studies. The labeling process involved pairwise comparisons of novel-view video sequences, with the resulting perceptual scores scaled into JOD units using Thurstone’s Case V psychometric model. This methodology produced continuous, scene-level perceptual quality estimates (Refer to [5] for a full description of the experimental design, scaling methodology, and subject demographics). It is important to highlight that the JOD scale’s utility and the comparative analysis it facilitates are particularly pronounced when conducted within the context of individual scenes [5].
@@
Evaluation protocol. We largely adhere to the evaluation protocol from [29, 30, 16], but in a more stringent manner. Our evaluation consists of two primary objectives: first, to determine if the proposed method can simultaneously adapt to three different datasets; second, to test the method’s ability to perform cross-dataset validation. Following the self-supervised learning phase, the model’s weights are frozen, and the model is applied to the three evaluation datasets to generate quality representations. A one-time linear regression is performed on a randomly sampled half of the integrated dataset to align these quality representations with human scores. The model’s performance is then evaluated on the remaining half of the dataset to assess the generalizability of the NVS-SQA framework across the integrated datasets. This approach is designed to thoroughly evaluate the framework’s capability to adapt and perform effectively amidst the varied complexities posed by the Lab, LLFF, and Fieldwork datasets simultaneously. Additionally, we conduct cross-dataset experiments to further assess the adaptability of the proposed method across various datasets. Importantly, the Lab and Fieldwork datasets were created and labeled by Liang et al.  [5] and are publicly available. The LLFF dataset constructed by Mildenhall et al. [10] is also publicly accessible. All three datasets were constructed by independent research groups prior to our work. This ensures that the cross-dataset evaluation reflects realistic generalization to external and independently produced NSS data.
@@
Training setup. The model was trained using the ADAM optimizer [64] for 200 epochs with a batch size of 16, resulting in a total training time of approximately 56.99 hours. Once training on the unlabeled dataset is complete, the model can be efficiently applied to novel scenes across varied datasets using only minor linear regression. The computational experiments were conducted on a high-specification desktop equipped with an AMD 5950X processor, an RTX 4090 GPU, and 128 GB of RAM, running on Windows 10 operating system. The implementation was carried out in PyTorch [65].
@@
TV +$0.378$ ($0.64$) +$0.423$ ($0.58$) +$0.311$ ($0.62$) +$0.087$ ($0.65$) +$0.050$ ($0.61$) +$0.075$ ($0.52$) +$0.200$ ($0.48$) +$0.136$ ($0.30$) $\underline{0.229}$ ($0.36$)
BRISQUE +$0.089$ ($0.58$) +$0.152$ ($0.58$) +$0.067$ ($0.53$) 0.037-0.037 ($0.54$) 0.103-0.103 ($0.45$) 0.050-0.050 ($0.44$) $\underline{0.214}$ ($0.83$) +$0.204$ ($0.72$) +$0.171$ ($0.70$)
NIQE +$0.467$ ($0.63$) +$0.331$ ($0.66$) +$0.400$ ($0.56$) +$0.025$ ($0.50$) 0.077-0.077 ($0.43$) +$0.025$ ($0.42$) 0.357-0.357 ($0.16$) 0.329-0.329 ($0.12$) 0.314-0.314 ($0.17$)
PIQE +$0.079$ ($0.49$) 0.079-0.079 ($0.55$) +$0.273$ ($0.57$) +$0.047$ ($0.56$) 0.079-0.079 ($0.47$) 0.021-0.021 ($0.44$) 0.364-0.364 ($0.65$) 0.060-0.060 ($0.70$) 0.302-0.302 ($0.49$)
CLIP-IQA +$0.233$ ($0.63$) +$0.178$ ($0.57$) +$0.200$ ($0.54$) +$0.025$ ($0.47$) 0.046-0.046 ($0.36$) +$0.025$ ($0.34$) 0.057-0.057 ($0.50$) 0.240-0.240 ($0.47$) 0.086-0.086 ($0.36$)
CONTRIQUE $\underline{0.689}$ ($0.28$) $\underline{0.759}$ ($0.29$) $\underline{0.622}$ ($0.29$) +$0.350$ ($0.40$) +$0.400$ ($0.50$) +$0.275$ ($0.33$) +$0.086$ ($0.43$) +$0.200$ ($0.53$) +$0.057$ ($0.36$)
Re-IQA +$0.589$ ($0.53$) +$0.585$ ($0.48$) +$0.489$ ($0.44$) +$0.062$ ($0.71$) 0.018-0.018 ($0.71$) +$0.025$ ($0.61$) +$0.143$ ($0.16$) $\underline{0.213}$ ($0.30$) +$0.200$ ($0.10$)
VIIDEO +$0.022$ ($0.37$) +$0.070$ ($0.44$) +$0.022$ ($0.28$) 0.050-0.050 ($0.49$) +$0.002$ ($0.49$) 0.000-0.000 ($0.39$) +$0.000$ ($0.27$) +$0.061$ ($0.47$) 0.029-0.029 ($0.30$)
Video-BlIINDS +$0.189$ ($0.39$) +$0.148$ ($0.41$) +$0.156$ ($0.31$) +$0.162$ ($0.44$) +$0.051$ ($0.40$) +$0.175$ ($0.38$) 0.314-0.314 ($0.33$) 0.113-0.113 ($0.42$) 0.286-0.286 ($0.33$)
FAST-VQA +$0.167$ ($0.59$) +$0.106$ ($0.59$) +$0.122$ ($0.46$) +$0.112$ ($0.55$) +$0.255$ ($0.55$) +$0.185$ ($0.41$) 0.171-0.171 ($0.65$) 0.191-0.191 ($0.63$) 0.186-0.186 ($0.52$)
FasterVQA +$0.186$ ($0.51$) +$0.245$ ($0.55$) +$0.172$ ($0.47$) +$0.162$ ($0.58$) +$0.176$ ($0.49$) +$0.125$ ($0.48$) 0.129-0.129 ($0.46$) 0.175-0.175 ($0.44$) 0.186-0.186 ($0.38$)
DOVER +$0.200$ ($0.56$) +$0.267$ ($0.65$) +$0.267$ ($0.50$) +$0.150$ ($0.61$) +$0.153$ ($0.47$) +$0.125$ ($0.48$) 0.129-0.129 ($0.26$) 0.209-0.209 ($0.18$) 0.214-0.214 ($0.22$)
DOVER-Mobile +$0.344$ ($0.60$) +$0.341$ ($0.61$) +$0.311$ ($0.47$) +$0.200$ ($0.47$) +$0.254$ ($0.36$) +$0.125$ ($0.40$) 0.071-0.071 ($0.47$) 0.097-0.097 ($0.32$) 0.143-0.143 ($0.46$)
NR-LFQA +$0.207$ ($0.52$) +$0.157$ ($0.56$) +$0.149$ ($0.42$) +$0.088$ ($0.47$) +$0.106$ ($0.34$) +$0.114$ ($0.37$) 0.094-0.094 ($0.59$) 0.067-0.067 ($0.57$) 0.039-0.039 ($0.51$)
Tensor-NLFQ +$0.282$ ($0.42$) +$0.290$ ($0.49$) +$0.202$ ($0.43$) +$0.223$ ($0.38$) +$0.210$ ($0.36$) +$0.125$ ($0.29$) 0.061-0.061 ($0.64$) 0.064-0.064 ($0.57$) 0.072-0.072 ($0.55$)
@@
LFACon +$0.333$ ($0.44$) +$0.310$ ($0.47$) +$0.242$ ($0.33$) $\underline{0.412}$ ($0.48$) $\underline{0.395}$ ($0.31$) $\underline{0.325}$ ($0.36$) +$0.157$ ($0.60$) +$0.132$ ($0.53$) +$0.086$ ($0.51$)
Proposed $\boldsymbol{0.911}$ ($0.08$) $\boldsymbol{0.883}$ ($0.11$) $\boldsymbol{0.822}$ ($0.16$) $\boldsymbol{0.700}$ ($0.28$) $\boldsymbol{0.644}$ ($0.39$) $\boldsymbol{0.625}$ ($0.19$) $\boldsymbol{0.700}$ ($0.04$) $\boldsymbol{0.678}$ ($0.10$) $\boldsymbol{0.571}$ ($0.09$)
@@
Figure 7: How often does each quality estimation method align with ground-truth perceptual labels (indicated by /’-shaped lines)? We compare our proposed NVS-SQA against three competitive representatives of no-reference quality assessment: CONTRIQUE (IQA), DOVER (VQA), and LFACon (LFIQA), evaluated on the Fieldwork, LLFF, and Lab datasets. Each subfigure shows scatter plots where different symbols and colors represent individual scenes, with a linear regression line for each. Each subfigure reports the number of scenes within the dataset where the predicted scores exhibit a positive correlation (positive slope) with human perceptual scores. For example, ’9/9’ in Fieldwork indicates that NVS-SQA achieved positive alignment in all 9 individual scenes.
@@ IV-B Comparison with no-reference quality assessment methods
Visualiazation of correlation with perceptual labels. Figure 7 illustrates how well the predicted quality assessments correlate with ground-truth perceptual labels. The ground-truth perceptual labels used in this analysis are the JOD scores described in Section III-C. These labels reflect human-perceived quality differences and were computed from pairwise comparison data as part of the public datasets in [5]. Each subplot corresponds to a dataset, and each regression line represents one individual scene within that dataset. The count (e.g., 9/9 for Fieldwork) reflects the number of scenes for which the model’s predicted quality scores positively correlate with human labels. This provides a scene-level consistency measure beyond global correlation coefficients. As shown in the figure, NVS-SQA achieves the highest number of positively sloped lines in every dataset, reaching perfect alignment in Fieldwork (9 of 9 lines) and top performance in both LLFF (7 of 8) and Lab (7 of 7). This result underscores the robust effectiveness of NVS-SQA across a diverse set of scenes and highlights its superior alignment with human perception compared to existing approaches.
@@
TV +$0.344$ ($0.58$) +$0.398$ ($0.58$) +$0.244$ ($0.54$) 0.050-0.050 ($0.50$) 0.032-0.032 ($0.53$) +$0.025$ ($0.42$) +$0.143$ ($0.51$) +$0.266$ ($0.27$) +$0.171$ ($0.46$)
BRISQUE 0.044-0.044 ($0.59$) 0.004-0.004 ($0.58$) 0.067-0.067 ($0.51$) 0.237-0.237 ($0.50$) 0.289-0.289 ($0.52$) 0.175-0.175 ($0.42$) $\underline{0.214}$ ($0.42$) +$0.279$ ($0.43$) +$0.171$ ($0.35$)
NIQE +$0.344$ ($0.37$) +$0.337$ ($0.40$) +$0.311$ ($0.33$) 0.212-0.212 ($0.58$) 0.138-0.138 ($0.55$) 0.175-0.175 ($0.50$) 0.243-0.243 ($0.22$) 0.188-0.188 ($0.41$) 0.171-0.171 ($0.22$)
PIQE 0.092-0.092 ($0.51$) 0.012-0.012 ($0.58$) 0.211-0.211 ($0.56$) 0.175-0.175 ($0.55$) 0.173-0.173 ($0.51$) 0.171-0.171 ($0.54$) 0.169-0.169 ($0.30$) 0.237-0.237 ($0.25$) 0.295-0.295 ($0.21$)
CLIP-IQA +$0.444$ ($0.34$) +$0.525$ ($0.32$) +$0.356$ ($0.25$) +$0.175$ ($0.60$) +$0.264$ ($0.67$) +$0.100$ ($0.49$) 0.057-0.057 ($0.25$) +$0.020$ ($0.26$) 0.029-0.029 ($0.22$)
CONTRIQUE $\underline{0.622}$ ($0.38$) $\underline{0.626}$ ($0.31$) $\underline{0.522}$ ($0.42$) 0.013-0.013 ($0.34$) +$0.050$ ($0.48$) 0.075-0.075 ($0.20$) 0.143-0.143 ($0.33$) 0.102-0.102 ($0.56$) 0.114-0.114 ($0.32$)